# A Trifecta of Modelling Tools: Navigating COMPAS, Shark and Bayesian Inference for Binary Black Hole Model Selection

### v1

@@ Abstract
Gravitational waves (GWs) have revealed surprising properties of binary black hole (BBH) populations, but there is still mystery surrounding how these compact objects evolve. We apply Bayesian inference and an efficient method to calculate the BBH merger rates in the Shark host galaxies, to determine the combination of COMPAS parameters that outputs a population most like the GW sources from the LVK transient catalogue. For our COMPAS models, we calculate the likelihood with and without the dependence on the predicted number of BBH merger events. We find strong correlations between hyper-parameters governing the specific angular momentum (AM) of mass lost during mass transfer, the mass-loss rates of Wolf–Rayet stars via winds and the chemically homogeneous evolution (CHE) formation channel. For the AM loss, we also find opposing trends in the likelihood with and without the dependence on the number of GW events. We conclude that analysing the marginalised and unmarginalised likelihood is a good indicator of whether the population parameters distribution and number of observed events reflect the LVK data. In doing so, we see that the majority of the models preferred in terms of the population-level parameters of the BBHs greatly overpredict the number of events we should have observed to date. Looking at the smaller number of models which perform well with both likelihoods, we find that those with no CHE, AM loss occurring closer to the donor during the first mass-transfer event, and/or higher rates of mass-loss from Wolf–Rayet winds are generally preferred by current data. We find these conclusions to be robust to our choice of selection criteria. However, we do identify several notable events that all of our models struggle to reproduce, and that, even when incorporating selection effects into our models to match our selection of LVK data, COMPAS tends to underpredict the proportion of events with chirp mass $>50M_{\odot}$ and $z>0.5$, leaving avenues for future investigations.
@@
pagerange: A Trifecta of Modelling Tools: Navigating COMPAS, Shark and Bayesian Inference for Binary Black Hole Model SelectionC
@@ 1 Introduction
    Isolated: The formation of merging BBHs through the evolution of isolated binaries. These stars are gravitationally bound since birth and are not perturbed by other stars or compact objects.
@@
    Dynamical: BBH mergers occurring in dense environments via three-body encounters followed by energy exchange and hardening of the binary. Some examples of dense environments are young star, nuclear or globular clusters, or in accretion disks of the central supermassive BHs in galaxies.
@@
GWs can be used to gain insight into the poorly constrained aspects of massive stellar binary evolution and help in distinguishing between binary formation channels. The detailed origins and formation channels for binary systems are still unknown, particularly what fraction of these systems come from isolated and dynamical channels.
@@
There are also aspects of stellar evolution that are poorly understood, such as the common envelope phase, mass transfer efficiency and angular momentum loss. This information is critical for understanding the population of BBHs observed by LIGO, Virgo and KAGRA (LVK), including their mass and redshift distributions, and to inform us on what is required for future detectors (Mapelli, 2019, 2020; Mandel & Broekgaarden, 2022; Arca Sedda et al., 2023). Combining the data from GWTC-1 (Abbott et al., 2019a), GWTC-2 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b), there is a total of 90 confident ($p_{\rm astro}>0.5$) GW event candidates observed to date, including 85 BBHs, 2 binary neutron stars (BNSs), and 3 black-hole-neutron-stars (BHNSs). This growing population of GWs, including those to be released from the ongoing fourth LVK observing run (O4), provides key information on the source masses, spins and local merger rates (Abbott et al., 2023a) of binary compact objects, which can be used to answer some of these questions.
@@
However, there are many formation pathways that could lead to the same individual source properties. Simulations of ensembles of events, produced by varying the physical assumptions, can provide better insight into how these binary systems evolve and merge over time. When combined with the full set of data observed to date, they can potentially break the degeneracy between uncertainties in stellar evolution parameters and BBH populations. They also enable us to produce merger rates that can be compared to observations. Mandel & Broekgaarden (2021) provide a comprehensive review of the local merger rates for the isolated and dynamical formation channels. The studies in the review combine population synthesis codes with different assumptions for the metallicity dependent star formation rate (SFR) density, which plays a crucial role in the evolution of massive binaries (Neijssel et al., 2019; Chruślińska, 2024). Broekgaarden et al. (2022a) utilises COMPAS (Stevenson et al., 2017; Vigna-Gómez et al., 2018; Team COMPAS: Riley, J. et al., 2022) to explore the impact of stellar evolution parameters and metallicity specific star formation, $\mathcal{S}(z,Z)$, on the merger rate, chirp mass and coalescence time distributions. They found that uncertainties in the binary evolution models only led to variations in the BBH merger rate of up to a factor 2, with the BBH merger rates most sensitive to $\mathcal{S}(z,Z)$. Santoliquido et al. (2021) and Neijssel et al. (2019) also come to similar conclusions. Many of these studies choose to explore the variations in binary evolution models over a grid of metallicities, and generate populations for each grid point. Some studies have shown that population III stars can also contribute to the total BBH merger rate density (Santoliquido et al., 2024; Tanikawa, 2024). These studies emphasise the need to further study the assumptions in binary evolution models to understand the impact on the merger rate, specifically the combined impact of model variations (Stevenson & Clarke, 2022).
@@
There have also been various studies proposing methods for constraining different binary evolution models. Wong et al. (2023) applies “backward" modelling of each GW event to the progenitors to avoid fixed astrophysical assumptions, by allowing variations in the hyper-parameters. However, they note that their methods become less efficient as the dimensionality of the parameter space increases. Barrett et al. (2018) uses Fisher information matrices to quantitatively analyse how the observed BBH population will constrain the COMPAS models, given the chirp mass distribution and accounting for the metallicity specific cosmic SFR, selection effects and measurement uncertainty. Riley & Mandel (2023) uses artificial neural networks to generate an interpolant to calculate the merger detection rates for double compact objects for varying prescriptions of the metallicity specific SFR. This is used to create GW populations for each model, where the chirp mass and redshift distribution are compared using Bayesian analysis to the LVK data. Delfavero et al. (2023) calculates a joint likelihood for binary evolution models. They fit normal distribution fits to the likelihood grids and implement Gaussian process regression to interpolate between models, the detected number of events and the likelihood. This is also used to quantify the systematic uncertainty in the parameters. However, they emphasise that the inference on a limited subspace of models could be biased. These studies highlight the high computational costs for multi-dimensional parameter studies. However, as more observations are made, the uncertain astrophysics will produce correlated impacts in the binary population, so there is a crucial need for more parameters covering a large range in model inference. Realistically, all formation channels (within the isolated and dynamical channels) must be considered to obtain a parameter space representative of the true BBH population (Zevin et al., 2021; Cheng et al., 2023a)
@@
To build on this body of work, we develop here a new method of combining a more accurate, non-parametric star-formation and metallicity history of the Universe with a wider range of different massive binary evolution models. We then use these models alongside data from GWTC-1 (Abbott et al., 2019a), GWTC-2 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b) to identify the most likely avenues for generating binary compact objects, and to rule out those that are unable to produce the distributions of objects seen to date. We do this by improving on Rauf et al. (2023), where we combined runs from the COMPAS rapid population synthesis code, for formation and evolution of BBH formed through isolated binary evolution, with a star formation rate and metallicity history for the observed host galaxies predicted by the semi-analytic galaxy evolution code Shark (Lagos et al., 2018; Lagos et al., 2019). In our original work, we predicted (among other things) the BBH merger rate as a function of redshift for four COMPAS runs, where we varied the remnant mass prescription. We found the merger rate ranged over several magnitudes and we concluded that the Mandel & Müller (2020) remnant mass prescription was the best fit out of the four models to the local merger rate currently predicted using the BBH mergers from GWTC-3. However, we also found that cross-matching between the Shark galaxies and the COMPAS population was prohibitively time-consuming for exploring larger portions of our parameter space. Since we explored only rates, we were also not able to confidently rule out any models that do not align with the GW population-level parameters, and did not test the impact of varying sub-parameters of these models such as those covering other aspects of binary evolution rather than just the mass of the remnant compact objects.
@@
So, in this paper, we substantially extend on our work in Rauf et al. (2023) and explore an alternative and efficient method of predicting the BBH merger rate and population level parameters such as the binary masses and redshift distributions. We do so by rescaling a high-fidelity single run of our original simulation, varying some of the most relevant COMPAS input parameters. This allows us to test in this work $>300$ different models for population synthesis, varying ingredients/model assumptions such as the common envelope efficiency, mass transfer accretion efficiency, mass-loss rates of Wolf–Rayet stars, chemically homogeneous evolution model and specific angular momentum loss. For each of these we evaluate the ability of the model to reproduce the observed distribution of black hole masses and redshifts, using a Bayesian framework to account for selection effects. We also explore the impact on our findings with and without including the merger rate as a fitting parameter (i.e., including a comparison of the model-predicted and observed merger rates in evaluating our model posteriors). Compared to previous studies this work also has the improvement of incorporating a basic estimate of sampling error in the simulated model merger rates and posteriors. This allows us to set a threshold for what we consider to be tangible differences in model performance, and also demonstrates more generally the importance of further investigation into the uncertainties in the typical ‘simulation-based inference’ used in GW population studies.
@@
This paper is structured as follows. In Section 2, we summarise recent progress in our understanding of various massive binary evolution models and parameters, motivating the range of models we explore in this work. In Section 3, we show that starting with a fiducial model, where the merger rate in a set of simulated galaxies has been evaluated by carefully cross-matching between COMPAS and Shark, we can rescale the simulation to produce the time and metallicity dependent GW merger rate for another model where the COMPAS parameters have been altered. In Section 4, we implement Bayesian inference to determine which of these COMPAS runs output a BBH population that best matches the current observed population. We do so by comparing the chirp mass, mass ratio and redshift distributions from each model with the LVK posterior sample distributions. We then validate the total number of theoretical and detected GW events expected for the models and the models posteriors by comparing to the values obtained using the four ‘brute-force’ evaluations from Rauf et al. (2023). We repeat the method for a wide range of new COMPAS runs in Section 5, identifying the best and worst fit models to the data and studying the impact of different selection effects and analysis choices on the conclusions we can draw about stellar evolution. We conclude in Section 6. In this work, we use the Shark cosmology $H_{0},\Omega_{m}=\{67.8,0.308\}$ where necessary.
@@ 2 COMPAS: Overview of Models
In the following section, we detail the physics and the implementation in COMPAS of the common envelope, mass transfer, mass loss via winds, angular momentum loss and formation channels such as the isolated and chemically homogeneous evolution. We review the literature on the assumptions and parameters associated with these processes that can impact the evolution of the binary stars and, hence, the number and source properties of GWs for each COMPAS model. We conclude with a summary of the big questions surrounding these prescriptions and their limitations in the context of BBH population modelling.
@@ 2.1 Isolated formation channels
In this paper, we focus on the isolated formation channel. This is the most commonly studied channel. It involves massive stars formed on the zero-age-main-sequence (ZAMS). We describe, one of the formation channels, the classical isolated binary evolution via a common envelope phase as follows:
@@
    The primary star in the binary reaches the end of main sequence (MS) evolution. Hydrogen burning in the core causes the outer stellar envelope to expand past the Roche lobe and transfer mass to the secondary star.
@@
    Mass transfer from the primary stops and leaves a Wolf-Rayet star. This then reaches the end of its life, undergoes a supernova and forms a compact object. The system is now an X-ray binary.
@@
    The secondary evolves and transfers mass back to the primary (compact) object. This process is dynamically unstable and forms a common envelope, which encompasses the compact object.
@@
    Orbital energy released due to the compact object rapidly in-spiralling can remove the common envelope.
@@
    The secondary star evolves and eventually also forms a compact object.
@@
    The double compact object (DCO) system slowly in-spirals due to orbital energy loss via GWs.
@@
However, there are many stages within this formation channel where the binary may not survive, or the events can subtly differ from the above prescription leading to dramatically different outcomes. The system may not survive the first supernova explosion or the common envelope may be not be ejected resulting in the compact object merging with the secondary core before a DCO is formed. Alternatively, the second mass transfer could be stable and avoid the common envelope (e.g., van den Heuvel et al., 2017; Neijssel et al., 2019; van Son et al., 2022a; Gallegos-Garcia et al., 2021; Marchant et al., 2021). The wind mass-loss rates and remnant prescription can also impact the orbital separation, as well as the final masses. There are various circumstances and factors that can affect the separation of DCO systems, which is strongly correlated with the delay time and the merger rate. On one hand, this makes GW sources unique probes of cosmic history (Neijssel et al., 2019; Mapelli, 2021; Mandel & Farmer, 2022; Costa et al., 2023), but on the other, leads to difficulties in interpreting or identifying the most important/critical points in the evolution of the system.
@@
As an example, Stevenson et al. (2017) used COMPAS to show that the first three GW events detected by LIGO: GW150914, GW151226 and GW151012, can be explained by the classical isolated channel with mass transfer and common envelope phase, and characterised their progenitor metallicity and masses. However, Dorozsmai & Toonen (2022) found that the stable channel, consisting of two stable mass transfer phases, dominates the GW sources — formation of GW sources via a radiative common envelope channel is not possible for low mass transfer efficiency ($f_{\rm MT}\leq 0.3$) due to the mass ratio being below the critical mass ratio ($q_{\rm crit}=M_{d}/M_{a}$, where $M_{a}$ is the acretor mass and $M_{d}$ is the donor mass). Picco et al. (2023) similarly found that stable mass transfer, where mass and angular momentum is lost through the L2 point, is a possible formation channel for BBHs. Comparing the two channels, van Son et al. (2022a) found that the common envelope channel tends to dominate BBH mergers with primary masses $<30M_{\odot}$ binaries with short delay times ($<1$Gyr). They also found that the common envelope phase is rare for massive stars due to heavy mass loss preventing them from filling their Roche lobe. As such, the stable mass transfer channel, which widens the binary orbit as mass is lost from the system during secondary mass transfer, forms primary BH masses $>30M_{\odot}$ and systems with long delay times ($>1$Gyr).

### v2

# A Trifecta of Modelling Tools: A Bayesian Binary Black Hole Model Selection combining Population Synthesis and Galaxy Formation Models

@@ Abstract
Gravitational waves (GWs) have revealed surprising properties of binary black hole (BBH) populations, but there is still mystery surrounding how these compact objects evolve. We apply Bayesian inference and an efficient method to calculate the BBH merger rates in the Shark host galaxies, to determine the combination of COMPAS parameters that outputs a population most like the GW sources from the LVK transient catalogue. For our COMPAS models, we calculate the likelihood with and without the dependence on the predicted number of BBH merger events. We find strong correlations between hyper-parameters governing the specific angular momentum (AM) of mass lost during mass transfer, the mass-loss rates of Wolf–Rayet stars via winds and the chemically homogeneous evolution (CHE) formation channel. We conclude that analysing the marginalised and unmarginalised likelihood is a good indicator of whether the population parameters distribution and number of observed events reflect the LVK data. In doing so, we see that the majority of the models preferred in terms of the population-level parameters of the BBHs greatly overpredict the number of events we should have observed to date. Looking at the smaller number of models which perform well with both likelihoods, we find that those with no CHE, AM loss occurring closer to the donor during the first mass-transfer event, and/or higher rates of mass-loss from Wolf–Rayet winds are generally preferred by current data. We find these conclusions to be robust to our choice of selection criteria.
@@
pagerange: A Trifecta of Modelling Tools: A Bayesian Binary Black Hole Model Selection combining Population Synthesis and Galaxy Formation ModelsC
@@ 1 Introduction
    Isolated: The formation of merging BBHs through the evolution of isolated binaries. These stars are gravitationally bound since birth and are not perturbed by other stars or compact objects (Mapelli & Giacobbo, 2018; Bavera et al., 2021; Santoliquido et al., 2021; Broekgaarden et al., 2021; Broekgaarden et al., 2022b; Olejak et al., 2022).
@@
    Dynamical: BBH mergers occurring in dense environments via three-body encounters followed by energy exchange and hardening of the binary. Some examples of dense environments are young star, nuclear or globular clusters, or in accretion disks of the central supermassive BHs in galaxies (Rodriguez et al., 2019; Yang et al., 2019; Arca Sedda et al., 2020; Zevin et al., 2021b).
@@
GWs can be used to gain insight into the poorly constrained aspects of massive stellar binary evolution and help in distinguishing between binary formation channels. The detailed origins and formation channels for binary systems are still unknown, particularly what fraction of these systems come from isolated, which is the focus of this paper, and dynamical channels.
@@
There are also aspects of stellar evolution that are poorly understood, such as the common envelope phase, mass transfer efficiency and AM loss. This information is critical for understanding the population of BBHs observed by LIGO, Virgo and KAGRA (LVK), including their mass and redshift distributions, and to inform us on what is required for future detectors (Mapelli, 2019, 2020; Mandel & Broekgaarden, 2022; Arca Sedda et al., 2023). Combining the data from GWTC-1 (Abbott et al., 2019a), GWTC-2 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b), there is a total of 90 confident ($p_{\rm astro}>0.5$; see Section 4.1 and Eq. 22) GW event candidates observed to date, including 85 BBHs, 2 binary neutron stars (BNSs), and 3 black-hole-neutron-stars (BHNSs). This growing population of GWs, including those to be released from the ongoing fourth LVK observing run (O4), provides key information on the source masses, spins and local merger rates (Abbott et al., 2023a) of binary compact objects, which can be used to answer some of these questions.
@@
However, there are many formation pathways that could lead to the same individual source properties. Simulations of ensembles of events, produced by varying the physical assumptions, can provide better insight into how these binary systems evolve and merge over time. When combined with the full set of data observed to date, they can potentially break the degeneracy between uncertainties in stellar evolution parameters and BBH populations. They also enable us to produce merger rates that can be compared to observations. Mandel & Broekgaarden (2021) provide a comprehensive review of the local merger rates for the isolated and dynamical formation channels. The studies in the review combine population synthesis codes with different assumptions for the metallicity dependent star formation rate (SFR) density, which plays a crucial role in the evolution of massive binaries (Neijssel et al., 2019; Chruślińska, 2024). Broekgaarden et al. (2022a) utilises COMPAS (Stevenson et al., 2017; Vigna-Gómez et al., 2018; Team COMPAS: Riley, J. et al., 2022) to explore the impact of stellar evolution parameters and metallicity-specific star formation, $\mathcal{S}(z,Z)$, on the merger rate, chirp mass and coalescence time distributions. They find that uncertainties in the binary evolution models only led to variations in the BBH merger rate of up to a factor 2, with the BBH merger rates most sensitive to $\mathcal{S}(z,Z)$. Santoliquido et al. (2021) and Neijssel et al. (2019) also come to similar conclusions. Mennekens & Vanbeveren (2014); Santoliquido et al. (2021); Chu et al. (2022) find variations in the merger rate over several orders of magnitude. Many of these studies choose to explore the variations in binary evolution models over a grid of metallicities, and generate populations for each grid point. Some studies have shown that population III stars can also contribute to the total BBH merger rate density (Santoliquido et al., 2024; Tanikawa, 2024). These studies emphasise the need to further study the assumptions in binary evolution models to understand the impact on the merger rate, specifically the combined impact of model variations (Stevenson & Clarke, 2022).
@@
There have also been various studies proposing methods for constraining different binary evolution models. Wong et al. (2023) applies “backward" modelling of each GW event to the progenitors to avoid fixed astrophysical assumptions, by allowing variations in the hyper-parameters. However, they note that their methods become less efficient as the dimensionality of the parameter space increases. Barrett et al. (2018) uses Fisher information matrices to quantitatively analyse how the observed BBH population will constrain the COMPAS models, given the chirp mass distribution and accounting for the metallicity-specific cosmic SFR, selection effects and measurement uncertainty. Riley & Mandel (2023) uses artificial neural networks to generate an interpolant to calculate the merger detection rates for double compact objects for varying prescriptions of the metallicity-specific SFR. This is used to create GW populations for each model, where the chirp mass and redshift distribution are compared using Bayesian analysis to the LVK data. Delfavero et al. (2023) calculates a joint likelihood for binary evolution models. They fit normal distribution fits to the likelihood grids and implement Gaussian process regression to interpolate between models, the detected number of events and the likelihood. This is also used to quantify the systematic uncertainty in the parameters. However, they emphasise that the inference on a limited subspace of models could be biased. These studies highlight the high computational costs for multi-dimensional parameter studies. However, as more observations are made, the uncertain astrophysics will produce correlated impacts in the binary population, so there is a crucial need for more parameters covering a large range in model inference. Realistically, all formation channels (within the isolated and dynamical channels) must be considered to obtain a parameter space representative of the true BBH population (Zevin et al., 2021a; Cheng et al., 2023).
@@
To build on this body of work, we develop here a new method of combining a more accurate, non-parametric star-formation and metallicity history of the Universe with a wider range of different massive binary evolution models. We then use these models alongside data from GWTC-1 (Abbott et al., 2019a), GWTC-2 (Abbott et al., 2021a) and GWTC-3 (Abbott et al., 2023b) to identify the most likely avenues for generating binary compact objects, and to rule out those that are unable to produce the distributions of objects seen to date. We do this by improving on Rauf et al. (2023), where we combined runs from the COMPAS rapid population synthesis code, for formation and evolution of BBH formed through isolated binary evolution, with a SFR and metallicity history for the observed host galaxies predicted by the semi-analytic galaxy evolution code Shark (Lagos et al., 2018; Lagos et al., 2019). In our original work, we predicted (among other things) the BBH merger rate as a function of redshift for four COMPAS runs, where we varied the remnant mass prescription. We found the merger rate ranged over several magnitudes and we concluded that the Mandel & Müller (2020) remnant mass prescription was the best fit out of the four models to the local merger rate currently predicted using the BBH mergers from GWTC-3. However, we also found that cross-matching between the Shark galaxies and the COMPAS population was prohibitively time-consuming for exploring larger portions of our parameter space. Since we explored only rates, we were also not able to confidently rule out any models that do not align with the GW population-level parameters, and did not test the impact of varying sub-parameters of these models such as those covering other aspects of binary evolution rather than just the mass of the remnant compact objects.
@@
So, in this paper, we substantially extend on our work in Rauf et al. (2023) and explore an alternative and efficient method of predicting the BBH merger rate and population level parameters such as the binary masses and redshift distributions. We do so by rescaling a high-fidelity single run of our original simulation, varying some of the most relevant COMPAS input parameters. This allows us to test in this work $>300$ different models for population synthesis, varying ingredients/model assumptions such as the common envelope efficiency, mass transfer accretion efficiency, mass-loss rates of Wolf–Rayet stars, chemically homogeneous evolution model and specific AM loss. For each of these we evaluate the ability of the model to reproduce the observed distribution of black hole masses and redshifts, using a Bayesian framework to account for selection effects. We also explore the impact on our findings with and without including the merger rate as a fitting parameter (i.e., including a comparison of the model-predicted and observed merger rates in evaluating our model posteriors). Compared to previous studies this work also has the improvement of incorporating a basic estimate of sampling error in the simulated model merger rates and posteriors. This allows us to set a threshold for what we consider to be tangible differences in model performance, and also demonstrates more generally the importance of further investigation into the uncertainties in the typical ‘simulation-based inference’ used in GW population studies.
@@
This paper is structured as follows. In Section 2, we summarise recent progress in our understanding of various massive binary evolution models and parameters, motivating the range of models we explore in this work. In Section 3, we show that starting with a fiducial model, where the merger rate in a set of simulated galaxies has been evaluated by carefully cross-matching between COMPAS and Shark, we can rescale the simulation to produce the time and metallicity dependent GW merger rate for another model where the COMPAS parameters have been altered. In Section 4, we implement Bayesian inference to determine which of these COMPAS runs output a BBH population that best matches the current observed population. We do so by comparing the chirp mass, mass ratio and redshift distributions from each model with the LVK posterior sample distributions. We then validate the total number of theoretical and detected GW events expected for the models and the models posteriors by comparing to the values obtained using the four ‘brute-force’ evaluations from Rauf et al. (2023). We repeat the method for a wide range of new COMPAS runs in Section 5, identifying the best and worst fit models to the data and studying the impact of different selection effects and analysis choices on the conclusions we can draw about stellar evolution. We conclude in Section 6. In this work, we use the Shark cosmology $H_{0},\Omega_{m}=\{67.8,0.308\}$ in flat $\Lambda$CDM where necessary.
@@ 2 COMPAS: Overview of Models
In the following section, we detail the physics and the implementation in COMPAS of the common envelope (CE), mass transfer, mass loss via winds, angular momentum (AM) loss and formation channels such as the isolated and chemically homogeneous evolution. We review the literature on the assumptions and parameters associated with these processes that can impact the evolution of the binary stars and, hence, the number and source properties of GWs for each COMPAS model. We conclude with a summary of the big questions surrounding these prescriptions and their limitations in the context of BBH population modelling. Note that in this section, we refer to the more massive star as the donor and the lighter star as the accretor star.
@@ 2.1 Isolated formation channels
In this paper, we focus on BBHs in the isolated formation channel. This is the most commonly studied channel, but see Fig 1 in Han et al. (2020) for possible binary formation channels. It involves massive stars formed on the zero-age-main-sequence (ZAMS). We describe, one of the formation channels, the classical isolated binary evolution via a CE phase as follows:
@@
    The primary star in the binary reaches the end of main sequence (MS) evolution. After hydrogen burning in the core ends, the core contracts and causes the outer stellar envelope to expand past the Roche lobe (due to the negative heat capacity of the star (Lynden-Bell & Wood, 1968)) and transfer mass to the secondary star.
@@
    Mass transfer from the primary stops and leaves a Wolf-Rayet star. [^1] This formation scenario for Wolf–Rayet stars is disputed by Shenar et al. (2020) due to lack of observations of massive stars at low metallicity. This then reaches the end of its life, undergoes a supernova and forms a BH. The system is now an X-ray binary.
@@
    The secondary evolves and transfers mass back to the primary BH. This process is dynamically unstable and forms a CE, which encompasses the compact object.
@@
    Orbital energy released due to the compact object rapidly in-spiralling can remove the CE.
@@
    The secondary star evolves and eventually also forms a BH.
@@
    The BBH system slowly in-spirals due to orbital energy loss via GWs.
@@
However, there are many stages within this formation channel where the binary may not survive, or the events can subtly differ from the above prescription leading to dramatically different outcomes. The system may not survive the first supernova explosion or the CE may be not be ejected resulting in the compact object merging with the secondary core before a BBH is formed (Schrøder et al., 2020; Grichener, 2023). Alternatively, the second mass transfer could be stable and avoid the CE (e.g., van den Heuvel, 1976; van den Heuvel et al., 2017; Neijssel et al., 2019; van Son et al., 2022a; Gallegos-Garcia et al., 2021; Marchant et al., 2021). The wind mass-loss rates and remnant prescription can also impact the orbital separation, as well as the final masses. There are various circumstances and factors that can affect the separation of BBH systems, which are strongly correlated with the delay time and the merger rate. On one hand, this makes GW sources unique probes of cosmic history (Neijssel et al., 2019; Mapelli, 2021; Mandel & Farmer, 2022; Costa et al., 2023), but on the other, leads to difficulties in interpreting or identifying the most important/critical points in the evolution of the system.
@@
As an example, Stevenson et al. (2017) used COMPAS to show that the first three GW events detected by LIGO: GW150914, GW151226 and GW151012, can be explained by the classical isolated channel with mass transfer and CE phase, and characterised their progenitor metallicity and masses. However, Dorozsmai & Toonen (2022) find that the stable channel, which consists of two stable mass transfer phases, dominates the GW sources. The formation of GW sources via a radiative CE channel is not possible for low mass transfer efficiency ($f_{\rm MT}\leq 0.3$, see Eq. 1 in Section 2.1.1) due to the mass ratio being below the critical mass ratio ($q_{\rm crit}=M_{d}/M_{a}$, where $M_{a}$ is the accretor mass and $M_{d}$ is the donor mass) for instability. Picco et al. (2023) similarly find that stable mass transfer, where mass and AM is lost through the L2 point, is a possible formation channel for BBHs.
Comparing the two channels, van Son et al. (2022a) find that the CE channel tends to dominate BBH mergers with primary masses $<30M_{\odot}$ in binaries with short delay times ($<1$Gyr). They also find that the CE phase is rare for massive stars due to heavy mass loss preventing them from filling their Roche lobe. As such, the stable mass transfer channel, which widens the binary orbit as mass is lost from the system during secondary mass transfer, forms primary BH masses $>30M_{\odot}$ and systems with long delay times ($>1$Gyr). They also find that the CE or stable mass transfer channel dominates the mass distribution depending on redshift, which is in line with Neijssel et al. (2019) finding redshift dependence in the mass distribution. van Son et al. (2022b) studies the stable mass transfer channel only and the impact of $f_{\rm MT}$ on the local BBH merger rate and the primary BH mass distribution. For $f_{\rm MT}=0.75$ and $f_{\rm MT}=1$, there are no BBHs formed through the stable mass transfer channel. The merger rate drops from $141~{}\rm Gpc^{-3}yr^{-1}$ with $f_{\rm MT}=0.25$ to $16~{}\rm Gpc^{-3}yr^{-1}$ with $f_{\rm MT}=0.5$, which implies the stable mass transfer contributes strongly to the merger rate for $f_{\rm MT}<0.5$. The low mass end of the primary BH mass distribution is justified by the stable mass transfer channel — as $f_{\rm MT}$ increases, the minimum primary BH mass increases. Overall, the relative contribution of both channels is uncertain. The CE channel in particular may be over-represented in population synthesis models (see Section 2.1.1 for relevant references).